Input: number of days on vacation, Output: total cost of vacation in dollars, Rule: Cost is $1590 plus $200 per day. Use integer inputs from 1 to 5. write the rule as an equation.
Answer by drj(1380)

Input: number of days on vacation, Output: total cost of vacation in dollars, Rule: Cost is $1590 plus $200 per day. Use integer inputs from 1 to 5. write the rule as an equation.

Step 1. Let x = number of days of vacation

Step 2. Let y=1590+200x be the total cost of vacation in dollars

Step 3. So the cost for various values of x consists of the following

x=1 y=1590+200(1)=$1790

x=2 y=1590+200(2)=$1990

x=3 y=1590+200(3)=$2190

x=2 y=1590+200(4)=$2390

x=2 y=1590+200(5)=$2590

I hope the above steps were helpful.

For free Step-By-Step Videos on Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra or for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

And good luck in your studies!

Respectfully,
Dr J